### v1

# Model Editing by Pure Fine-Tuning

Abstract
Fine-tuning is dismissed as not effective for model editing due to its poor performance compared to more specialized methods. However, fine-tuning is simple, agnostic to the architectural details of the model being edited, and able to leverage ongoing advances in standard training methods (e.g., PEFT), making it an appealing choice for a model editor. In this work, we show that pure fine-tuning can be a viable approach to model editing. We propose a slight modification of naive fine-tuning with two key ingredients. First, we optimize the conditional likelihood rather than the full likelihood. Second, we augment the data with random paraphrases and facts to encourage generalization and locality. Our experiments on ZsRE and CounterFact show that this simple modification allows fine-tuning to often match or outperform specialized editors in the edit score.
Model Editing by Pure Fine-Tuning
Govind Gangadhar and Karl Stratos Department of Computer Science Rutgers University {govind.gangadhar, karl.stratos}@rutgers.edu
1 Introduction
Model editing is a promising approach to combating incorrect or otherwise unwanted knowledge in LLMs. It assumes a set of edits that assert desired information and aims to alter the model so that it not only succeeds in memorizing the edits (efficacy), but also applying the asserted information to new prompts (generalization), without affecting inference that should remain unchanged (locality). There is clearly a trade-off between these metrics. At one extreme, the model can achieve high efficacy by memorizing the edits, but it will fail in generalization and locality. At another extreme, the model can achieve high locality by staying unmodified, but it will fail in efficacy and generalization.
A straightforward approach of fine-tuning the model on the requested edits is well known to perform poorly, especially in locality. This motivated researchers to dismiss fine-tuning as an ineffective solution to model editing and develop very specialized methods such as MEND (Mitchell et al., 2022), ROME (Meng et al., 2022), and MEMIT (Meng et al., 2023). A central theme of these methods is “minimally invasive modification”, typically involving a careful selection of weights and rank-one updates. While technically interesting, they require a suite of assumptions which may not be satisfied in other contexts (e.g., with a different model architecture). In contrast, fine-tuning is simple, completely agnostic to the details of the model, and can take advantage of advances in standard training such as parameter-efficient fine-tuning (PEFT) techniques.
Editor
Standard
model?
Batched
edits?
No
training?
Effective
edits?
Naive fine-tuning
MEND
ROME
MEMIT
Our fine-tuning
Table 1: Conceptual comparisons of model editors.
In this work, we show that fine-tuning can be a viable approach to model editing, focusing primarily on improving the edit score. We shift the focus from models and algorithms to training objectives and data augmentation. See Table 1 for conceptual comparisons between our approach and other editors. Our method is a slight variation on naive fine-tuning with two small but important differences. First, in line with the theme of minimal modification, we optimize the conditional likelihood (i.e., mask all tokens except the edited target). Second, we augment the data with self-induced paraphrases of the prompt and arbitrary unrelated facts to encourage generalization and locality. This simple change is sufficient for pure fine-tuning to match or outperform specialized editors in mass-editing and also perform respectably in single-editing.
2 Task
We consider the task of mass-editing. Let $\mathcal{V}$ denote the vocabulary. A fact is a sentence $x\in\mathcal{V}^{T}$ that expresses a subject-relation-object triple $(s,r,o)$ in natural language. We follow the convention in the model editing literature and assume that the object $o\in\mathcal{V}^{m}$ is the last $m$ tokens of $x$. The prefix $\pi=(x_{1}\ldots x_{m-1})$ expresses $(s,r)$ and is denoted as the prompt. Let $\mathcal{E}$ denote a set of facts to enforce (i.e., requested edits). Our goal is to edit a pretrained language model so that it upholds the relations expressed in $\mathcal{E}$ without changing its behavior on other facts. In ZsRE, $\mathcal{E}$ consists of 10,000 factual statements (e.g., “The artwork Gideon’s Way was by who? John Creasey”). In CounterFact, $\mathcal{E}$ consists of 10,000 counterfactual statements (e.g., “TextEdit, a product of Nintendo”).
An editor is evaluated by three competing metrics: efficacy, generalization, and locality. Let $p_{\theta}$ denote a language model edited on $\mathcal{E}$. In ZsRE, efficacy is the accuracy of $o=\operatorname*{arg\,max}_{y}p_{\theta}(y|\pi)$ for $(\pi,o)\in\mathcal{E}$; generalization is the accuracy of $o=\operatorname*{arg\,max}_{y}p_{\theta}(y|\pi_{\text{par}})$ where $\pi_{\text{par}}$ is some paraphrase of $\pi$; locality is the accuracy of $o_{\text{unrel}}=\operatorname*{arg\,max}_{y}p_{\theta}(y|\pi_{\text{unrel}})$ where $(\pi_{\text{unrel}},o_{\text{unrel}})$ is an unrelated fact. In CounterFact, efficacy and generalization measure the accuracy of $p_{\theta}(o|\pi)>p_{\theta}(o_{\text{true}}|\pi)$ and $p_{\theta}(o|\pi_{\text{par}})>p_{\theta}(o_{\text{true}}|\pi_{\text{par}})$ where $o_{\text{true}}$ is a true object for $(\pi,o)\in\mathcal{E}$ (e.g., “Apple” in “TextEdit, a product of Nintendo”). Locality measures the accuracy of $p_{\theta}(o_{\text{true}}|\pi_{\text{nb}})>p_{\theta}(o|\pi_{\text{nb}})$ where $\pi_{\text{nb}}$ is a “neighbhorhood” prompt whose target must remain unchanged from $o_{\text{true}}$ (e.g., “Macintosh File System, a product of” $\rightarrow$ “Apple”). In both datasets, the final edit score is the harmonic mean of efficacy, generalization, and locality.
It is well known that naive fine-tuning, namely just optimizing
$\min_{\theta}\;-\sum_{x\in\mathcal{E}}\;\log p_{\theta}(x)$ (1)
results in a poor editing score. The main reason is that while it improves efficacy and possibly even generalization, it harms locality by changing the model’s predictions on unrelated or neighborhood prompts. This led to the development of highly specialized editing methods, with the implicit assumption that fine-tuning is not effective for model editing.
3 Method
We propose a purely fine-tuning-based method that achieves a competitive editing score. Our objective is a slight variation of (1):
$\min_{\theta}\;-\sum_{(\pi,o)\in\mathcal{E}\cup\mathcal{P}\cup\mathcal{R}}\;\log p_{\theta}(o|\pi)$ (2)
There are two small but important differences between (1) and (2). First, we optimize the conditional likelihood of the edit target $o|\pi$ rather than the full likelihood. Our motivation is to make the training more focused (for efficacy and generalization) and minimize the damage caused by fine-tuning (for locality).
Second, we fine-tune not only on the requested edits $\mathcal{E}$ but also additional facts in $\mathcal{P}$ and $\mathcal{N}$ to promote generalization and locality. $\mathcal{P}$ is pseudo-paraphrases of the prompts in $\mathcal{E}$ generated by appending text generated by the (unedited) model. More specifically, for each $(\pi,o)\in\mathcal{E}$ we generate $((z,\pi),o)\in\mathcal{P}$ where $z\sim p_{\theta}$. $\mathcal{R}$ is a set of facts that should not be altered by editing on $\mathcal{E}$. There are many ways to obtain $\mathcal{R}$, but we find taking random facts from the training split to be simple and effective. We filter $x\in\mathcal{R}$ to ensure that $x$ does not include the subject-relation-object triple used in evaluation.
We emphasize that the augmented sets $\mathcal{P}$ and $\mathcal{R}$ do not use the evaluation facts (otherwise it is trivial). Given the requested edits $\mathcal{E}$, we perform the data augmentation and finetune the model according to (2).
Contrastive learning.
Given that our goal in CounterFact is to achieve $p_{\theta}(o|\pi)>p_{\theta}(o_{\text{true}}|\pi)$ where $o$ is the new object to enforce over the original $o_{\text{true}}$ for a given prompt $\pi$, it is natural to consider a contrastive objective. We experimented with DPO (Rafailov et al., 2023) where we frame $o|\pi$ as the “preferred” response over $o_{\text{true}}|\pi$, optimizing the DPO loss jointly with (2). However, we found that prompt masking and data augmentation are already effective and do not benefit from contrastive learning.
ZsRE CounterFact
Editor Score Efficacy Generalization Locality Score Efficacy Generalization Locality
— (original GPT-J) 26.4 26.4 (0.6) 25.8 (0.5) 27.0 (0.5) 22.4 15.2 (0.7) 17.7 (0.6) 83.5 (0.5)
FT-W (21st layer w/ weight decay) 42.1 69.6 (0.6) 64.8 (0.6) 24.1 (0.6) 67.6 99.4 (0.1) 77.0 (0.7) 46.9 (0.6)
MEND 20.0 19.4 (0.5) 18.6 (0.5) 22.4 (0.5) 23.1 15.7 (0.7) 18.5 (0.7) 83.0 (0.5)
ROME 2.6 21.0 (0.7) 19.6 (0.7) 0.9 (0.1) 50.3 50.2 (1.0) 50.4 (0.8) 50.2 (0.6)
MEMIT 50.7 96.7 (0.3) 89.7 (0.5) 26.6 (0.5) 85.8 98.9 (0.2) 88.6 (0.5) 73.7 (0.5)
FT 44.8 99.9 (0.03) 98.9 (0.2) 21.4 (0.5) 52.8 79.6 (0.8) 58.5 (0.8) 36.8 (0.7)
FT (21st layer) 42.9 99.9 (0.0) 87.4 (0.5) 20.5 (0.5) 60.5 99.9 (0.04) 63.3 (0.8) 42.0 (0.6)
FT + Mask 58.3 97.6 (0.3) 91.7 (0.5) 32.9 (0.6) 54.3 97.1 (0.3) 62.1 (0.8) 34.7 (0.6)
FT + Mask + Para 56.1 99.9 (0.0) 98.7 (0.2) 29.9 (0.5) 63.7 100.0 (0.0) 92.5 (0.4) 38.0 (0.6)
FT + Mask + Para + Rand 62.0 99.9 (0.0) 97.0 (0.3) 35.6 (0.6) 86.5 98.8 (0.2) 93.6 (0.4) 72.0 (0.6)
FT + Mask + Para + Rand + DPO 85.5 98.8 (0.2) 93.4 (0.4) 70.1 (0.6)
Table 2: Mass-editing results on ZsRE and CounterFact (10,000 edits each) with GPT-J. The results of the specialized methods (FT-W, MEND, ROME, MEMIT) are from Meng et al. (2023). FT denotes naive fine-tuning on the requested edits; FT (21st layer) denotes FT only on the 21st layer. “+ Mask” means we mask the prompt. “+ Para” means we augment the data with model-generated paraphrases of the requested edits. “+ Rand” means we augment the data with random facts from the training split (not overlapping with any evaluation facts). “+ DPO” means we additionally optimize the DPO loss term using the changed target as the preferred response over the original target (which is supplied in CounterFact).
4 Related Work
We discuss some prominent model editing methods and the different assumptions they require to highlight the simplicity of our fine-tuning approach; we refer to Zhang et al. (2024) for an in-depth survey.
MEND (Mitchell et al., 2022) is a meta-learning method (Sinitsin et al., 2020; De Cao et al., 2021) that predicts the change in the gradient. Similar to our approach, it also involves an explicit training stage (on the training split of ZsRE). It uses a special rank-one decomposition of the gradient for parameter efficiency. In our case, we achieve parameter efficiency without special considerations simply by leveraging PEFT methods. MEND works poorly for mass-editing.
ROME (Meng et al., 2022) is a locate-then-edit method (Geva et al., 2021; Dai et al., 2022) that first uses causal tracing to identify the feedforward layer to change, and then applies a rank-one update to the layer’s weight matrix. While it does not involve an explicit training stage, it does require an explicit knowledge of the subject $s$ in the edit and its vector representation (which is heuristically induced). It also relies on Wikipedia to estimate the covariance matrix of the subject embeddings (needed in the rank-one update). ROME works well on single-editing but poorly on mass-editing. MEMIT (Meng et al., 2023) extends ROME to multiple edits by carefully spreading updates to multiple layers. Like ROME, it requires layer specification, subject embeddings, and covariance statistics obtained from Wikipedia. MEMIT is very effective for mass-editing and is our strongest baseline.
IKE (Zheng et al., 2023) proposes in-context learning for model editing. While it similarly uses demonstrations, it is critically limited to single-editing. It also requires an initial model that is capable of effective in-context learning. Consequently, IKE focuses on single-editing with large-scale LLMs, making their results not comparable to ours.
We define our scope as achieving a competitive edit score through pure fine-tuning. In particular, we do not focus on preserving general capabilities of the model being edited. Existing methods consider basic metrics such as $n$-gram repetition, but recent work shows that performance on a variety of downstream tasks drops to zero under any editor (Gu et al., 2024). We leave preserving the general capabilities of an LLM with a fine-tuning editor as the next step of our work.
ZsRE CounterFact
Editor Score Efficacy Generalization Locality Score Efficacy Generalization Locality
— (original GPT-2 XL) 22.5 22.2 (0.5) 21.3 (0.5) 24.2 (0.5) 30.5 22.2 (0.9) 24.7 (0.8) 78.1 (0.6)
FT 45.9 99.6 (0.1) 82.1 (0.1) 23.2 (0.5) 65.1 100.0 (0.0) 87.9 (0.6) 40.4 (0.7)
FT-L 40.1 92.3 (0.4) 47.2 (0.7) 23.4 (0.5) 66.9 99.1 (0.2) 48.7 (1.0) 70.3 (0.7)
KN - - - - 35.6 28.7 (1.0) 28.0 (0.9) 72.9 (0.7)
KE 41.8 65.5 (0.6) 61.4 (0.6) 24.9 (0.5) 52.2 84.3 (0.8) 75.4 (0.8) 30.9 (0.7)
MEND 42.9 75.9 (0.5) 65.3 (0.6) 24.1 (0.5) 57.9 99.1 (0.2) 65.4 (0.9) 37.9 (0.7)
ROME 47.9 99.8 (0.0) 88.1 (0.5) 24.2 (0.5) 89.2 100 (0.1) 96.4 (0.3) 75.4 (0.7)
FT + Mask + Para + Sim 51.4 100.0 (0.0) 99.9 (0.0) 26.1 (0.8) 83.1 98.6 (0.3) 87.3 (0.7) 69.0 (0.7)
Table 3: Single-editing results on ZsRE (10,000 edits) and CounterFact (7,500 edits) with GPT-2 XL. “+ Sim” means we include similar facts (measured by Sentence-BERT) instead of random facts for locality supervision.
Editor Score Fluency Consistency
— (original GPT-J) 22.4 622.4 (0.3) 29.4 (0.2)
FT-W 67.6 293.9 (2.4) 15.9 (0.3)
MEND 20.0 618.4 (0.3) 31.1 (0.2)
ROME 50.3 589.6 (0.5) 3.3 (0.0)
MEMIT 85.8 619.9 (0.3) 40.1 (0.2)
FT 52.8 626.1 (0.4) 31.0 (0.2)
FT + Mask 54.3 563.6 (0.5) 6.1 (0.1)
FT + Mask + Para 63.7 550.7 (0.6) 4.7 (0.1)
FT + Mask + Para + Rand 86.5 352.0 (1.5) 5.2 (0.2)
Table 4: Fluency and consistency with mass-editing on CounterFact.
5 Experiments
5.1 Mass-Editing
Our main results are on mass-editing with on ZsRE and CounterFact, following the same setting in Meng et al. (2023). Each dataset provides 10,000 requested edits (i.e., $\mathcal{E}$). As described in Section 3, we augment the fine-tuning data with pseudo-paraphrases of the edit prompts $\mathcal{P}$ for generalization supervision and random facts $\mathcal{R}$ for locality supervision. More specifically, for each edit in $\mathcal{E}$ we generate 15 paraphrases and take 20 facts from the training split while ensuring that they do not contain any evaluation facts. Thus the total number of facts we fine-tune on is 360,000. We fine-tune GPT-J (6B) (Wang and Komatsuzaki, 2021) with LoRA (Hu et al., 2022) for computational efficiency. We optimize (2). The training takes around 7 hours on 8 GPUs.
We give the results in Table 2 where the row “FT + Mask + Para + Rand” corresponds to our final method. We show various ablations to illustrate the effect of individual components. We can make the following observations. First, the proposed prompt masking in (2) (FT + Mask) improves the performance of vanilla fine-tuning (FT), supporting the intuition that minimizing the model update is beneficial. In fact, on ZsRE this already outperforms MEMIT without any data augmentation. Second, augmenting the data with paraphrased prompts substantially improves generalization (FT + Mask + Para). Third, augmenting the data with random facts further improves locality (FT + Mask + Para + Rand), allowing pure fine-tuning to match the performance of MEMIT on CounterFact. We find that adding the DPO loss on top of the final method does not yield improvements.
5.2 Single-Editing
To compare with existing methods developed for single-editing (i.e., updating the model for a single fact), we consider optimizing (2) one edit at a time. We follow the setting in ROME and use GPT-2 XL (1.5B) Radford et al. (2019). The training takes around 1.4 seconds per edit on average. We give the results in Table 3. We find that it is helpful to include similar facts instead of random facts for locality supervision, presumably because fine-tuning is more sensitive with small data size thus requiring more care in the selection of demonstrations. We use Sentence-BERT embeddings (Reimers and Gurevych, 2019) and take 15 nearest facts as $\mathcal{R}$. On ZsRE, our method outperforms all existing single-edit methods. On CounterFact, our method lags behind ROME which is specifically optimized for single-editing, but still achieves the second-best edit score.
Editor Score Fluency Consistency
— (original GPT-2 XL) 30.5 626.6 (0.3) 31.9 (0.2)
FT 65.1 607.1 (1.1) 40.5 (0.3)
FT-L 66.9 621.4 (1.0) 37.4 (0.3)
KN 35.6 570.4 (2.3) 30.3 (0.3)
KE 52.2 586.6 (2.1) 31.2 (0.3)
MEND 57.9 624.2 (0.4) 34.8 (0.3)
ROME 89.2 621.9 (0.5) 41.9 (0.3)
FT + Mask + Para + Sim 83.1 557.0 (1.8) 15.2 (0.3)
Table 5: Fluency and consistency with single-editing on CounterFact.
5.3 Generative metrics
We report results on generative metrics in Table 4 and 5 for completeness. Following prior work (Meng et al., 2022, 2023), we report fluency (entropy of $n$-gram distributions) and consistency (similarity score with a reference text). Our fine-tuning methods take a hit on these metrics while improving the edit score, showing that more work is needed to go beyond classification. However, recent work shows that none of the compared editors preserves downstream performance (Gu et al., 2024), thus achieving this goal meaningfully is still an open problem.
6 Conclusions
We have demonstrated that pure fine-tuning is sufficient to obtain strong edit performance with a slight modification: optimizing the conditional likelihood and augmenting the data with additional facts. Our results challenge the assumption that fine-tuning is ineffective as a model editor, suggesting that model editing could become part of standard training.
Limitations
Our scope is deliberately limited to the classification aspect of model editing rather than the generative aspect to demonstrate the viability of a fine-tuning editor. Thus our approach is currently unable to preserve the LLM’s general capabilities, as is the case with all existing mass editors.